OpenStudy (anonymous):

myininaya? The slope of a railway is expressed as a percent and is called the grade of the road. At the Royal Gorge, in Colorado, an inclined railway takes visitors down to the arkansas river. If the grade is 175% and the vertical drop of the railway is 1015 feet, what is the horizontal change of the railway?

myininaya (myininaya):

ok so slope is rise/run \[1.75=\frac{rise}{run}=\frac{1015}{run}\]

myininaya (myininaya):

We need to solve \[1.75=\frac{1015}{run}\] for run

myininaya (myininaya):

run is the horizontal change

OpenStudy (anonymous):

okay and then?

OpenStudy (anonymous):

is that all??

myininaya (myininaya):

\[1.75 run=1015\]

myininaya (myininaya):

yes just divide both sides by 1.75

OpenStudy (anonymous):

and that's the answer??

myininaya (myininaya):

whatever 1015/1.75 is yes

OpenStudy (anonymous):

105?

OpenStudy (anonymous):

that is the horizontal change?

myininaya (myininaya):

yes

myininaya (myininaya):

i get 580

myininaya (myininaya):

1015/1.75 is what i did

OpenStudy (anonymous):

ohh careless mistake. Thank you!~